OpenStudy (zzr0ck3r):

\(J:C_1([0,1])\rightarrow C_1([0,1])\) defined by \(J(f)(x)=\int_0^xf(t)dt\) I need to find \(|||J|||\) I have the definition \(|||L|||:=\text{sup}\{||L(x)|| \ : \ ||x||\le 1\}\). where \(L\) is a linear operation. I am super confused on how to even start going about this.... does this not mean that \(|||J|||=\text{sup}\{||J(f)|| \ : \ ||f||\le 1\}\) but then what is \(||f||\)? does \(||f||=|||f|||\)?

OpenStudy (zzr0ck3r):

@eliassaab

OpenStudy (anonymous):

Do you mean [0,10] or [0,1]? Do you mean D(f) or J(f)?

OpenStudy (zzr0ck3r):

sorry edited

OpenStudy (zzr0ck3r):

Suppose \(\max_{x\in[0,1]}|f(x)|\le 1$, then $f(x)\le1\) for \(x\in[0,1]\). Thus \(\int_0^xf(t)dt\le\int_0^x1*dt=x\le 1\) for \(\forall \ x\in [0,1].\) So, \(\max_{x\in[0,1]}|J(f)|\le 1\) and thus \(|||J|||\le1\) Now let \(f(x)=1\), then \(f(x)\in C([0,1])\) and \(\max_{x\in[0,1]}f(x)=1\) and \(\max_{x\in[0,1]}|\int_0^x1*dt= 1\). So \(\text{sup}\{\max_{x\in[0,1]}|\int_0^xf(t)dt| \ : \ \max_{x\in[0,1]}|f(x)|\le 1\}\ge 1\). Thus we have that \(|||J|||=1\).

OpenStudy (zzr0ck3r):

this is what I came up with, but not sure if I am way off or onto something.

OpenStudy (zzr0ck3r):

its from a part in the book where it says its easy to show |||J|||=1. I hate when it says its easy, it never is.....\(\huge :(\)

OpenStudy (zzr0ck3r):

sorry about the $ $ just imagine they are not there:)

OpenStudy (anonymous):

You did it right.

OpenStudy (zzr0ck3r):

oh wow, good deal

OpenStudy (zzr0ck3r):

ty

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

There is a small technical adjustment since f can be <0, You should have said that \[ || J(f)||=\sup_{|x|\le 1}\left | \int_0^x f(t) dt \right|\le \\ \sup_{|x|\le 1}\int_0^x\left | f(t) \right|dt\le \int_0^1\left | f(t) \right|dt\le \int_0^1 \sup_{t\in [0,1]}| f(t)| dt=\\ \sup_{t\in [0,1]}| f(t)|=||f|| \]

OpenStudy (zzr0ck3r):

great thnx

OpenStudy (anonymous):

YW

OpenStudy (zzr0ck3r):

sorry im a little confused on where that would need to be, where did mine brake?

OpenStudy (anonymous):

if \[ \int_0^xf(t)dt\le 1\] this does not imply \[ \left|\int_0^xf(t)dt\right|\le 1\]

OpenStudy (anonymous):

It does if you follows my steps

OpenStudy (zzr0ck3r):

oh yeah I see, good deal

OpenStudy (zzr0ck3r):

So i have to show that the following linear functional is continuous on C([0,1]). L(f)=f(0). So ||f||<=1 implies f(x)<=1 for all x in [0,1] implies f(0)<=1 all x in [0,1], so |||L|||<=1 so L is continuous, to show |||L||| =1 i just need to let f=1 then max f =1 showing |||L|||=1. Is this one right?

OpenStudy (anonymous):

\[ |L(f)|=|f(0)| \le ||f|\implies |||L|||<1\\ L(1)=1\implies |||L|||=1 \]

OpenStudy (zzr0ck3r):

@eliassaab can you explain this step please \(\large\int_0^1 \sup_{t\in [0,1]}| f(t)| dt=\sup_{t\in [0,1]}| f(t)|\) TY in advance.

OpenStudy (anonymous):

Let \[ M=\sup_{t\in [0,1]}| f(t)| dt \] then \[ \int_0^1 M dt = M \int_0^1 dt = M(1)=M \]

OpenStudy (zzr0ck3r):

ahh sup|f(t)| is just a real number, i get it:)